Grade 12th passMarks and Rank in IIT JEE

If f:[-20,20] Is relation defined by f(x) =[x^2/a] since+cosx is even where [.] Is the g.i.f then set of values of a given by Options (a) fie (B) R (C) [0,400](D) [400, infinity]

Aditya Pathak
8 Years agoGrade 12th pass

1 Answer

Dhruv
8 Years ago
f(x) - f(-x) = 0 (even function) => [x^2/a] sinx + cosx - (-[x^/a]sinx + cosx)=> 2×[x^2/a] sinx =0Now for making [x^2/a] =0 a should be greater than 400 .Hence, a belongs to (400,infinity)